# Visual Instruction Bottleneck Tuning

### v1

@@ Abstract
Despite widespread adoption, multimodal large language models (MLLMs) suffer performance degradation when encountering unfamiliar queries under distribution shifts. Existing methods to improve MLLM generalization typically require either more instruction data or larger advanced model architectures, both of which incur non-trivial human labor or computational costs. In this work, we take an alternative approach to enhance the robustness of MLLMs under distribution shifts, from a representation learning perspective. Inspired by information bottleneck (IB) principle, we derive a variational lower bound of the IB for MLLMs and devise a practical implementation, Visual Instruction Bottleneck Tuning (Vittle). We then provide a theoretical justification of Vittle by revealing its connection to an information-theoretic robustness metric of MLLM. Empirical validation of three MLLMs on open-ended and closed-form question answering and object hallucination detection tasks over 45 datasets, including 30 shift scenarios, demonstrates that Vittle consistently improves the MLLM’s robustness under shifts by pursuing the learning of a minimal sufficient representation.
@@ 1 Introduction
In intensive races on the track of frontier-level AI models, we have observed unprecedented achievements through the form of a general-purpose chat assistant known as multimodal large language models (MLLMs) xai_grok3_2025 ; openai_o3_o4mini_2025 ; googlecloud_gemini2_5_pro_2025 that combine a visual encoder with a large language model. Their universal yet flexible question-answering interface enables MLLMs to easily permeate our lives from general problem-solving liang2024survey ; yang2024large to practical applications AlSaad2024vy ; li2024large ; caffagni2024revolution ; Cui_2024_WACV . While these models may achieve human-like or even surpass human-level performance on certain tasks, a critical gap remains in their robustness—particularly in handling input variations that humans process effortlessly.
@@
Specifically, MLLMs fail to produce relevant responses under query distribution shifts. That is, they are vulnerable to processing subtly perturbed samples and long-tail samples oh2025understanding . This limitation partially stems from the difficulty of acquiring diverse high-quality multimodal instruction data at scale. When trained via standard maximum likelihood estimation on this relatively limited amount of instruction data, MLLM tends to fit to data-specific patterns and result in a brittle solution geirhos2020shortcut ; ye2024mmspubench ; liang2025aligning . To enhance generalization, existing efforts typically fall into two categories (1) data-centric approaches, which collect more instruction data zhao2023svit ; li2024llava ; gu2024infinity and processes input in a finer granularity liu2024llavanext ; shen2025long , and (2) model-centric approaches, which scale up the underlying model using more expressive or specialized backbones chen2024expanding ; tong2024cambrian ; shi2025eagle ; bai2025qwen2 . However, both data scaling and model scaling are resource-intensive—requiring significant annotation or computational cost.
@@ 3.1 Preliminary: Information Bottleneck As a Learning Objective
In other words, the IB objective promotes representations that discard non-essential features tied to the input modality, while preserving those critical for solving the task. This property is particularly desirable for robust instruction tuning, where diverse multimodal inputs must be mapped to consistent, meaningful outputs under varied conditions (e.g., visual and textual perturbations). Despite its appeal, integrating the IB objective into MLLM training is highly non-trivial due to the intractability of mutual information estimation and the complexity of autoregressive and multimodal architectures.
@@ 3.2 Variational Inference for Information Bottleneck in MLLMs
Directly optimizing the IB objective is generally intractable, as it involves mutual information terms over unknown data distributions. In this work, we introduce a tractable variational bound on the IB objective, specifically tailored to the autoregressive and multimodal structure of MLLMs. We outline the key steps below and provide full derivations in the Appendix C.
@@
where the first inequality holds given the non-negativity of Kullback-Leibler divergence (KLD), $D_{\rm KL}(r(z)||p(z))$, and $p(z_{v}|x_{v},x_{t})=p(z_{v}|x_{v})$ due to causal attention in MLLM. We introduce $r(z)=r(z_{v},z_{t})=r(z_{v})r(z_{t})$ as a factorizable variational approximation of the true prior $p(z)$.
Next, for the output-relevant term $I(Z,Y)$, we have the lower bound:
@@
where we replace the true posterior $p(y|z)$ with a variational approximation $q(y|z)$ that will be parameterized by a model (will be elucidated in Section 3.3).
@@ 3.3 Vittle: A Practical Implementation of Visual Instruction Bottleneck Tuning
By using Monte Carlo approximation of expectations over data, Eq. (3.2) can be expressed as follows,
$\mathcal{L}_{\beta}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{z|x^{i}}[\log q(y^{i}|z)]-\beta\;(D_{\rm KL}(p(z_{v}|x^{i}_{v})||r(z_{v}))+D_{\rm KL}(p(z_{t}|x^{i}_{v},x^{i}_{t})||r(z_{t}))).$ (5)
To compute this empirical estimate of the IB lower bound, we need to model the posterior distributions, $p(z_{v}|x_{v})$ and $p(z_{t}|x_{v},x_{t})$, and prior distributions $r(z_{v})$ and $r(z_{t})$, of the MLLM’s inner representation $Z$. While in principle these distributions can take arbitrary forms, multivariate Gaussian distributions have been widely adopted in variational inference and probabilistic embedding literature graves2011practical ; KingmaWelling2014 ; Blei03042017 ; alemi2017deep ; oh2018modeling ; chun2021probabilistic due to their mathematical tractability and empirical effectiveness. By following this common standard, we set the posteriors and priors as Gaussian with diagonal covariance for $d$-dimensional variable, and will elucidate how exactly they are defined below.
@@
As illustrated in Figure 2, we parameterize the posteriors $p(z_{v}|x_{v})$ and $p(z_{t}|x_{v},x_{t})$ using simple MLP blocks. Specifically, we introduce two non-linear projections, $g_{\phi_{v}},g_{\phi_{t}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2d}$, which map each $d$-dimensional token embedding to the posterior Gaussian parameter vectors $\mu\in\mathbb{R}^{d}$ and $\sigma^{2}\in\mathbb{R}_{+}^{d}$ for the vision and language modalities, respectively. Given an intermediate $l$-th layer representation $(z_{v},z_{t})=f_{\theta^{l}}(x_{v},x_{t})$, we define:
$p(z_{v}|x_{v})=\mathcal{N}(z_{v};\mu_{v},\sigma_{v}^{2}\cdot I),\quad p(z_{t}|x_{v},x_{t})=\mathcal{N}(z_{t};\mu_{t},\sigma_{t}^{2}\cdot I),$
where $[\mu_{v},\sigma_{v}^{2}]=g_{\phi_{v}}(f_{\theta^{l}}(x_{v}))$ and $[\mu_{t},\sigma_{t}^{2}]=g_{\phi_{t}}(f_{\theta^{l}}(x_{v},x_{t}))$, with the mean and variance parameters split along output dimensions of MLP. These MLPs are applied position-wise in the same manner as Transformer’s feed-forward layers vaswani2017attention , producing token-wise variational posteriors. Now, we can sample from the posterior distributions of MLLM representation by $\tilde{z}_{v}\sim p(z_{v}|x_{v})$ and $\tilde{z}_{t}\sim p(z_{t}|x_{v},x_{t})$. Then, to strike a balance between invariance and sensitivity, we interpolate the original representation $z$ (pre-bottleneck) with its bottlenecked counterpart $\tilde{z}$ as $\hat{z}=(1-\alpha)z+\alpha\tilde{z}$. These representations are fed into the remaining layers to compute the predictive distribution over outputs, i.e., $q(y|z):=f_{\theta^{l+}}(y|\hat{z}_{v},\hat{z}_{t})$. While direct sampling introduces non-differentiability, we can enable the gradient flow using the reparameterization trick KingmaWelling2014 to sample $\tilde{z}$ via $\tilde{z}=\mu+\sigma\odot\epsilon$ with $\epsilon\sim\mathcal{N}(\mathbf{0},I)$ where $\mu$ and $\sigma$ are the outputs of the bottleneck MLP module given input $x$.
@@
We consider two instantiations of the prior distribution for both $Z_{v}$ and $Z_{t}$: (1) a fixed standard Gaussian $\mathcal{N}(\mathbf{0},I)$, which is input-independent and enforces strong isotropy, and (2) a learnable Gaussian $\mathcal{N}(\mu_{\psi},\sigma^{2}_{\psi}\cdot I)$, where $\mu_{\psi}$ and $\sigma^{2}_{\psi}$ are two learnable vectors shared across samples. Each prior affects the formation of representations differently—the fixed prior imposes stronger regularization and robustness, while the learnable prior introduces additional flexibility by allowing the model to adapt to the instruction tuning distribution. We name the former Vittle (F) and the latter Vittle (L), and validate them altogether for all the evaluations in Section 4.
@@
The first term of $\mathcal{L}_{\beta}($Eq. (5)) can be easily computed through the standard cross-entropy, and our Gaussian instantiation of posterior and priors allows us to derive closed-form expressions of KLD terms that can be computed from simple arithmetic between $\mu$ and $\sigma^{2}$ parameters (See Appendix A.2). We set $\beta=\frac{0.1}{d}$ where $d$ is the hidden dimension of the MLLM, to normalize the KL regularization terms relative to the size of the latent dimension. The interpolation coefficient $\alpha$ in $\hat{z}=(1-\alpha)z+\alpha\tilde{z}$ increases progressively following a cosine schedule up to $0.5$. During inference, we consistently use an averaged representation $\hat{z}=(z+\tilde{z})/2$. The target layer to apply the bottleneck module can differ between visual and textual tokens, but we set $l=24$ for both modalities among 32 layers in a 7B-size LLM, i.e., top 25% layer, by default for simplicity (See Appendix B.1 for the ablation study). Figure 2 depicts the architecture overview.
@@ 3.4 Theoretical Justification
The learning objective of Vittle has an attractive theoretical interpretation that can support the improvement in robustness of Vittle. In this section, we first introduce a recently proposed information-theoretic measure of MLLM’s robustness under distribution shifts, effective mutual information difference (EMID oh2025understanding ), and show how Vittle can contribute to improving EMID.
@@
where $I(\cdot)$ denotes mutual information that measures the relevance between input instruction and response. A higher value of EMID indicates that MLLM $P_{\Theta}$ undergoes performance degeneration in the distribution $Q$ (test data) compared to $P$ (training data), so we want to achieve a lower value of it to ensure robustness. We now derive an upper bound for EMID (See Appendix D for the proof).
@@
Let $P_{\Theta}$ be an MLLM that maps $X=\{X_{v},X_{t}\}$ to $Z=\{Z_{v},Z_{t}\}$, and then sequentially maps $Z$ to $Y_{\Theta}$. Given joint distributions $P_{XY}=P_{X}\times P_{Y|X}$ and $Q_{XY}=Q_{X}\times Q_{Y|X}$ (resp. $P_{ZY}$ and $Q_{ZY}$), by assuming consistent conditionals over $Z_{v}|Z_{t}$, $Z_{t}|Z_{v}$, and $Y|X$ between $P$ and $Q$, we have an upper bound for $\text{EMID}(P_{XY},Q_{XY};{P_{\Theta}})$ as below,
@@
where $H$ and $D^{\frac{1}{2}}_{\rm JS}$ indicate the entropy and square root of Jensen-Shannon divergence (JSD), respectively, $\Delta_{X|Z}:=\mathbb{E}_{z\sim P}[D_{\rm KL}(P_{X|z}||M_{X|z})]+\mathbb{E}_{z\sim Q}[D_{\rm KL}(Q_{X|z}||M_{X|z})]$ with a mixture distribution $M=\frac{P+Q}{2}$, and $\hat{H}:=\max_{x\in\mathcal{X}}[H(Q_{Y|x})+H(P_{Y_{\Theta}})]$. As we consider an optimization problem of $\Theta$, the terms, $H(P_{Y})$, $H(Q_{Y})$, and $\max H(Q_{Y|X})$, can be ignored from Eq. 7. We can also ignore $\sqrt{\Delta_{X|Z}}$ term because it cannot directly affect $Y_{\Theta}$ given the Markov assumption $X\rightarrow Z\rightarrow Y_{\Theta}$.
Implication. Vittle maximizes the variational lower bound of IB, which consists of (1) minimizing a standard negative log-likelihood term representing an expected risk, and (2) minimizing KLD terms to enforce posterior distributions close to prior distributions. By pursuing (1), MLLM $P_{\Theta}$ seeks a solution $\Theta$ that minimizes the expected risk and reduces its output entropy $H(P_{Y_{\Theta}})$ and $H(Q_{Y_{\Theta}})$ wen2024mitigating ; yang2024can ; groot2024overconfidence . Besides, it also reduces JSD between representation distributions $P_{Z}$ and $Q_{Z}$ by promoting all posterior samples to be laid near the pre-defined priors. In summary, reduced entropy and JSD terms induce lower EMID, which means that Vittle tries to achieve minimal difference between effective mutual information over training and evaluation distributions, while adapting to the in-distribution training set.
We show that Vittle indeed reduces JSD and EMID under distribution shifts in Table 4.2, and demonstrate in Section 4.2 that Vittle’s nice theoretical property is translated into consistent robustness gains under 30 distribution shift scenarios while maintaining in-distribution task performance.

### v2

@@ Abstract
Despite widespread adoption, multimodal large language models (MLLMs) suffer performance degradation when encountering unfamiliar queries under distribution shifts. Existing methods to improve MLLM generalization typically require either more instruction data or larger advanced model architectures, both of which incur non-trivial human labor or computational costs. In this work, we take an alternative approach to enhance the generalization and robustness of MLLMs under distribution shifts, from a representation learning perspective. Inspired by information bottleneck (IB) principle, we derive a variational lower bound of the IB for MLLMs and devise a practical implementation, Visual Instruction Bottleneck Tuning (Vittle). We then provide a theoretical justification of Vittle by revealing its connection to an information-theoretic robustness metric of MLLM. Empirical validation of multiple MLLMs on open-ended and closed-form question answering and object hallucination detection tasks over 45 datasets, including 30 shift scenarios, demonstrates that Vittle consistently improves the MLLM’s robustness under shifts by pursuing the learning of a minimal sufficient representation.
Code: https://github.com/deeplearning-wisc/vittle
@@ 1 Introduction
In intensive races on the track of frontier-level AI models, we have observed unprecedented achievements through the form of a general-purpose chat assistant known as multimodal large language models (MLLMs) openai2025gpt5 ; anthropic2025claude45 ; comanici2025gemini ; xai2025grok4 that combine a visual encoder with a large language model. Their universal yet flexible question-answering interface enables MLLMs to easily permeate our lives from general problem-solving liang2024survey ; yang2024large to practical applications AlSaad2024vy ; li2024large ; caffagni2024revolution ; Cui_2024_WACV . While these models may achieve human-like or even surpass human-level performance on certain tasks, a critical gap remains in their robustness—particularly in handling input variations that humans process effortlessly.
@@
Specifically, MLLMs fail to produce relevant responses under query distribution shifts. That is, they are vulnerable to processing subtly perturbed samples and long-tailed samples oh2025understanding . This limitation partially stems from the difficulty of acquiring diverse high-quality multimodal instruction data at scale. When trained using standard maximum likelihood estimation on this relatively limited amount of instruction data, MLLM tends to fit to data-specific patterns and results in a brittle solution geirhos2020shortcut ; ye2024mmspubench ; liang2025aligning . To enhance generalization, existing efforts typically fall into two categories (1) data-centric approaches, which collect more instruction data zhao2023svit ; li2024llava ; gu2024infinity and processes input in a finer granularity liu2024llavanext ; shen2025long , and (2) model-centric approaches, which scale up the underlying model using more expressive or specialized backbones chen2024expanding ; tong2024cambrian ; shi2025eagle ; bai2025qwen2 . However, both data scaling and model scaling are resource-intensive—requiring significant annotation or computational cost.
@@ 3.1 Preliminary: Information Bottleneck As a Learning Objective
In other words, the IB objective promotes representations that discard non-essential features tied to the input domain, while preserving those critical for solving the task. This property is desirable for robust instruction tuning, where user queries that have the same latent goal must be mapped to consistent responses under varied conditions (e.g., visual and textual perturbations). Despite its appeal, integrating the IB objective into MLLM training is highly non-trivial due to the intractability of mutual information estimation and the complexity of autoregressive and multimodal architectures.
@@ 3.2 Variational Inference for Information Bottleneck in MLLMs
Directly optimizing the IB objective is generally intractable, as it involves mutual information terms over unknown data distributions. In this work, we introduce a tractable variational bound on the IB objective, specifically tailored to the autoregressive and multimodal structure of MLLMs. We outline the key steps below and provide full derivations in the Appendix D.
@@
where the first inequality holds given the non-negativity of Kullback-Leibler divergence (KLD), $D_{\rm KL}(r(z)||p(z))$, and $p(z_{v}|x_{v},x_{t})=p(z_{v}|x_{v})$ due to causal mask in MLLM. Here, we introduce $r(z)=r(z_{v},z_{t})=r(z_{v})r(z_{t})$ as a factorizable variational approximation of the true prior $p(z)$.
Next, for the output-relevant term $I(Z,Y)$, we have the lower bound as the same as Alemi et al. (alemi2017deep, ):
@@
where we replace the true posterior $p(y|z)$ with a variational approximation $q(y|z)$ that will be parameterized by a model component (will be elucidated in Section 3.3).
@@ 3.3 Vittle: A Practical Implementation of Visual Instruction Bottleneck Tuning
By using a Monte Carlo approximation of expectations over data, Eq. (3.2) can be written as follows,
$\mathcal{L}_{\beta}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{z|x^{i}}[\log q(y^{i}|z)]-\beta\;(D_{\rm KL}(p(z_{v}|x^{i}_{v})||r(z_{v}))+D_{\rm KL}(p(z_{t}|x^{i}_{v},x^{i}_{t})||r(z_{t}))),$ (5)
where, $x^{i}=(x_{v}^{i},x_{v}^{t})$ denotes the $i$-th sample query from an instruction tuning dataset. To compute this empirical estimate of the IB lower bound, we need to model the posterior distributions, $p(z_{v}|x_{v})$ and $p(z_{t}|x_{v},x_{t})$, and prior distributions $r(z_{v})$ and $r(z_{t})$, of the MLLM’s inner representation $Z$. Although in principle these distributions can take arbitrary forms, multivariate Gaussian distributions with simplified covariance matrices have been widely adopted in variational inference and probabilistic embedding literature graves2011practical ; KingmaWelling2014 ; Blei03042017 ; alemi2017deep ; oh2018modeling ; chun2021probabilistic ; chun2024improved ; chun2025probabilistic due to their mathematical tractability and empirical effectiveness. By following this common standard, we set the posteriors and priors as Gaussian with diagonal covariance, and will elucidate how exactly they are defined below.
@@
As illustrated in Figure 2, we parameterize the posteriors $p(z_{v}|x_{v})$ and $p(z_{t}|x_{v},x_{t})$ using simple feed-forward blocks. Specifically, they are non-linear mappings, $g_{\phi_{\cdot}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2d}$ implemented by multi-layer perceptron (MLP) for each modalities, which map each $d$-dimensional token embedding to the posterior Gaussian parameter vectors $\mu\in\mathbb{R}^{d}$ and $\sigma^{2}\in\mathbb{R}_{+}^{d}$. Given an intermediate $l$-th layer output representation $(z_{v},z_{t})=f_{\theta^{l}}(x_{v},x_{t})$, we define:
$p(z_{v}|x_{v}):=\mathcal{N}(z_{v};\mu_{v},\sigma_{v}^{2}\cdot I),\quad p(z_{t}|x_{v},x_{t}):=\mathcal{N}(z_{t};\mu_{t},\sigma_{t}^{2}\cdot I),$
where $\mu_{v},\sigma_{v}^{2}]=g_{\phi_{v}}(f_{\theta^{l}}(x_{v}))$ and $\mu_{t},\sigma_{t}^{2}]=g_{\phi_{t}}(f_{\theta^{l}}(x_{v},x_{t}))$, with the mean and variance parameters are bipartited along the output dimensions of the MLP. These MLPs are applied position-wise in the same manner as Transformer’s feed-forward layers vaswani2017attention , producing token-wise variational posteriors. Now, we can sample from the posterior distributions of MLLM representation by $\tilde{z}_{v}\sim p(z_{v}|x_{v})$ and $\tilde{z}_{t}\sim p(z_{t}|x_{v},x_{t})$. Then, to strike a balance between invariance and sensitivity, we interpolate the original representation $z$ (pre-bottleneck) with the post-bottleneck counterpart $\tilde{z}$ as $\hat{z}=(1-\alpha)z+\alpha\tilde{z}$ [^2] This interpolation yields an asymmetric posterior specification for $I(Z,X)$ and $I(Z,Y)$ (Appendix D).. These representations are fed into the remaining layers to compute the predictive distribution over outputs, i.e., $q(y|z):=f_{\theta^{l+}}(y|\hat{z}_{v},\hat{z}_{t})$. While direct sampling introduces non-differentiability, we can enable the gradient flow using the reparameterization trick KingmaWelling2014 to sample $\tilde{z}$ via $\tilde{z}=\mu+\sigma\odot\epsilon$ with $\epsilon\sim\mathcal{N}(\mathbf{0},I)$ where $\mu$ and $\sigma$ are the outputs of the bottleneck MLP module given input $x$.
@@
We consider two instantiations of the prior distribution for both $Z_{v}$ and $Z_{t}$: (1) a fixed standard [^3] An alternative is to use the representation statistics (mean and variance) from a pre-instruction-tune model to host informative priors similar to mixout (Lee2020Mixout, ), which may relax the need for interpolation $(1-\alpha)z+\alpha\tilde{z}$. Gaussian $\mathcal{N}(\mathbf{0},I)$, which is input-independent and enforces strong isotropy, and (2) a learnable Gaussian $\mathcal{N}(\mu_{\psi},\sigma^{2}_{\psi}\cdot I)$, where $\mu_{\psi}$ and $\sigma^{2}_{\psi}$ are two learnable vectors shared across samples. Each prior affects the formation of representations differently—the fixed prior imposes stronger regularization and robustness, while the learnable prior introduces additional flexibility by allowing the model to adapt to the instruction tuning distribution. We name the former Vittle (F) and the latter Vittle (L), and validate them altogether for all the evaluations in Section 4.
@@
The first term of $\mathcal{L}_{\beta}($Eq. (5)) can be easily computed through the standard cross-entropy, and our Gaussian instantiation of posteriors and priors allows us to derive closed-form expressions of KLD terms that can be computed from simple arithmetic between $\mu$ and $\sigma^{2}$ parameters (See Appendix A.2). We set $\beta=\frac{0.1}{d}$ where $d$ is the hidden dimension of the MLLM, to normalize the KL regularization terms relative to the size of the latent dimension. The interpolation coefficient $\alpha$ in $\hat{z}=(1-\alpha)z+\alpha\tilde{z}$ increases progressively following a cosine schedule up to $0.5$. During inference, we consistently use an averaged representation $\hat{z}=(z+\tilde{z})/2$ with a deterministic posterior representation $\tilde{z}=\mu$ rather than using a random sample. The target layer to apply the bottleneck module can differ between visual and textual tokens, but we set $l=24$ for both modalities among 32 layers in a 7B-size LLM, i.e., top 25% layer, by default for simplicity (See Appendix B.1 for the ablation study). Figure 2 depicts the architecture overview.
@@ 3.4 Theoretical Justification
The learning objective of Vittle has an attractive theoretical interpretation that can support the improvement in robustness of Vittle. In this section, we first introduce a recently proposed information-theoretic measure of MLLM’s robustness under distribution shifts, effective mutual information difference, EMID oh2025understanding , and show how Vittle can contribute to reducing EMID.
@@
Let $P_{\Theta}:\mathcal{X}\rightarrow{\mathcal{Y}}$ be an MLLM with parameters $\Theta$ that produces an output response $Y_{\Theta}$ given an input instruction $X$. For joint distributions $P_{XY}$ and $Q_{XY}$, effective mutual information difference of $P_{\Theta}$ over distributions $P$ and $Q$ is defined as below,
@@
where $I$ denotes the mutual information that measures the relevance between input query and response. A higher value of EMID indicates that MLLM $P_{\Theta}$ undergoes query-response relavance degeneration in the distribution $Q$ (e.g., test-time) compared to $P$ (e.g., train-time), so we want to achieve a lower value of it to ensure robustness. We now derive an upper bound for the EMID (See Appendix E).
@@
Let $P_{\Theta}$ be an MLLM that maps $X=\{X_{v},X_{t}\}$ to $Z=\{Z_{v},Z_{t}\}$, and then subsequently maps $Z$ to $Y_{\Theta}$. Given joint distributions $P_{XY}=P_{X}\times P_{Y|X}$ and $Q_{XY}=Q_{X}\times Q_{Y|X}$ (resp. $P_{ZY}$ and $Q_{ZY}$), by assuming consistent conditionals over $Z_{v}|Z_{t}$, $Z_{t}|Z_{v}$, and $Y|X$ between $P$ and $Q$, we have an upper bound for $\text{EMID}(P_{XY},Q_{XY};{P_{\Theta}})$ as below,
@@
where $H$ and $D^{\frac{1}{2}}_{\rm JS}$ indicate the entropy and square root of Jensen-Shannon divergence (JSD), respectively, $\Delta_{X|Z}:=\mathbb{E}_{z\sim P}[D_{\rm KL}(P_{X|z}||M_{X|z})]+\mathbb{E}_{z\sim Q}[D_{\rm KL}(Q_{X|z}||M_{X|z})]$ with a mixture distribution $M=\frac{P+Q}{2}$, and $\hat{H}:=\max_{x\in\mathcal{X}}[H(Q_{Y|x})+H(P_{Y_{\Theta}})]$. As we are interested in the terms related to $\Theta$ being optimized, the terms $H(P_{Y})$, $H(Q_{Y})$, and $\max H(Q_{Y|X})$, can be ignored from Eq. 7 which are fixed across model parameters. We can also ignore $\sqrt{\Delta_{X|Z}}$ term because it cannot directly affect $Y_{\Theta}$ given the conditional independence $X\bot Y_{\Theta}|Z$. That is, the upper bound of EMID is boiled down to the multiplication and summation between the output entropy and the representational divergence.
Implication. Vittle maximizes the variational lower bound of IB, which consists of (1) minimizing a standard negative log-likelihood term representing an expected risk, and (2) minimizing KLD terms to enforce posterior distributions close to prior distributions. By (1), MLLM $P_{\Theta}$ seeks a solution $\Theta$ that minimizes the expected risk and reduces its output entropy $H(P_{Y_{\Theta}})$ and probably $H(Q_{Y_{\Theta}})$ as well wen2024mitigating ; yang2024can ; groot2024overconfidence . By (2), it may reduce JSD between representation distributions $P_{Z}$ and $Q_{Z}$ by promoting all posterior samples to be laid near the pre-defined priors alemi2018fixing ; fu2019cyclical ; wang2021posterior . In summary, reduced entropy and JSD terms induce a lower EMID, which means that Vittle may achieve better robustness to distribution shifts than the standard training method that neglects the divergence in representation space.
We show that Vittle indeed reduces empirical JSD and EMID under distribution shifts in Table 4.2, and demonstrate in Section 4.2 that Vittle’s nice theoretical property is translated into consistent robustness gains under 30 distribution shift scenarios while maintaining in-distribution performance.